Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
10.25
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.5
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
10.75

Beginnings of night, Overcast, skipped TIC67240043 (first object). Later "improved" (cleared( but all night long relative humidity was around 75% and {Tenv - DewPoint} around 4C. Both DIMM2 and Pachon seeing monitor down. Seeing values are estimates based in our guider camera. Last third of night RH dropped to 64% and {Tenv - DewPoint} grew to 6C.
Seeing values are estimates based in our guider camera. DIMM2 back around 8 UT
Skipped, no time for: HIP023516, HIP025119, 55383975, 55383975